10a + 4b = 28
6a + 18b = 48

Answer by Edwin McCravy(20056)


10a +  4b = 28
 6a + 18b = 48

Divide the first one through by 2
Divide the second one through by 6

5a + 2b = 14
 a + 3b =  8

Solve the second for a:

      a = 8 - 3b

Substitute (8 - 3b) for a in the first:

       5a + 2b = 14
5(8 - 3b) + 2b = 14
 40 - 15b + 2b = 14
      40 - 13b = 14
          -13b = -26
             b = 2

Substitute (2) for b in 

             a = 8 - 3b
             a = 8 - 3(2)
             a = 8 - 6
             a = 2

Edwin